1 Find the gen sol of the given higher-order DE. 7 you 5y+ 3yt 9y=0
Answer #1

Given Differential equation is y-syu By 4 ay =o. -5074 8849=0 Stol . : - 11 -5 39 & I -6 9 10 :0H) CDN-6D+9)=0 (0-6D+9 -0. D- General Solution is (4) 8= (C +GX ) 3x + czer y = (a + Caxle & en Gog and g are Constants.